Economy Unemployment

One of the economic problems that always haunt a country is unemployment. The following shows the population data of the State "A". Total Population 285; The number of children over the age of 16 and elderly 65 years of age is 75, not the workforce is 50 and according to the Central Bureau of Statistics unemployment is 16%. Determine the number of unemployed, the ratio of workers to population and the participation rate of the labor force.
Answer #1

1. The number of unemployed is 12.

2. Participation rate in the labor force is 0.84

3. Ratio of workers to population is \(22.11 \%\).

Explanation:

1. Unemployment is \(16 \%\) from 75 people, so tu calculate the number of unemloyed multiply \(16 \%\) with 75 \(16 \% \times 75=12\)

2. The labor force participation rate is a measure of an economy's active workforce. The formula for the number is the sum of all employed or actively seeking employment divided by the total noninstitutionalized. Divide the number of people in the labor force (63) by the total adult (working-age) population (75) \(63 \div 75=0.84\)

3. Ratio of workers to population, divide 67 with total population and muliply with 100 \(\frac{63}{285} \times 100=22.11 \%\)

answered by: ForlifeGod